### v1

# Scalable Bilevel Loss Balancing for Multi-Task Learning

Abstract
Multi-task learning (MTL) has been widely adopted for its ability to simultaneously learn multiple tasks. While existing gradient manipulation methods often yield more balanced solutions than simple scalarization-based approaches, they typically incur a significant computational overhead of $\mathcal{O}(K)$ in both time and memory, where $K$is the number of tasks. In this paper, we propose BiLB4MTL, a simple and scalable loss balancing approach for MTL, formulated from a novel bilevel optimization perspective. Our method incorporates three key components: (i) an initial loss normalization, (ii) a bilevel loss-balancing formulation, and (iii) a scalable first-order algorithm that requires only $\mathcal{O}(1)$ time and memory. Theoretically, we prove that BiLB4MTL guarantees convergence not only to a stationary point of the bilevel loss balancing problem but also to an $\epsilon$-accurate Pareto stationary point for all $K$ loss functions under mild conditions. Extensive experiments on diverse multi-task datasets demonstrate that BiLB4MTL achieves state-of-the-art performance in both accuracy and efficiency. Code is available at https://github.com/OptMN-Lab/-BiLB4MTL.
Figure 1: The loss trajectories of a toy 2-task learning problem from Liu et al. 2024 and the runtime comparison of different MTL methods for 50000 steps. Stars on the Pareto front denote the converge points. Although FAMO achieves more balanced results than LS and MGDA, it converges to different points on the Pareto front. Our method reaches the same balanced point with a computational cost comparable to the simple Linear Scalarization (LS). Full experimental details can be found in Section A.1.
1 Introduction
In recent years, Multi-Task Learning (MTL) has received increasing attention for its ability to predict multiple tasks simultaneously using a single model, thereby reducing computational overhead. This versatility has enabled a wide range of applications, including autonomous driving (Chen et al., 2018), recommendation systems (Wang et al., 2020), and natural language processing (Zhang et al., 2022).
Typically, research in MTL follows two main schemes. Scalarization-based methods, such as linear scalarization, reduce MTL to a scalar optimization problem by using an averaged or weighted sum of loss functions as the objective. Due to its simplicity and scalability, it became the prominent approach in the early studies (Caruana, 1997). However, it often causes performance degradation compared with single-task learning due to the gradient conflict (Yu et al., 2020; Liu et al., 2021a). Gradient conflict arises from two main reasons: 1) gradients point in different directions and 2) gradient magnitudes vary significantly. As a result, the final update gradient may either be offset or dominated by the largest gradient (Liu et al., 2021b). To mitigate this issue, various gradient manipulation methods have been developed to find balanced and fair solutions via seeking a better conflict-aware update direction (Désidéri, 2012; Liu et al., 2021a; Ban & Ji, 2024; Navon et al., 2022; Yu et al., 2020; Fernando et al., 2023; Xiao et al., 2024). Nevertheless, most of these methods require computing and storing the gradients for all $K$ tasks, resulting in substantial $O(K)$ time and memory costs. This limitation reduces their scalability in large-scale MTL applications involving complex models and extensive datasets.
In this paper, we propose a simple and scalable loss balancing approach for MTL from a novel bilevel optimization perspective. Our approach comprises three key components: initial loss normalization, a bilevel loss balancing formulation, and a scalable first-order algorithmic design. Our specific contributions are summarized as follows.
  • $\bullet$
    Bilevel loss balancing. At the core of our bilevel formulation, the lower-level problem optimizes the model parameters by minimizing a weighted sum of normalized individual loss functions. Meanwhile, the upper-level problem adjusts these weights to minimize the disparities among the loss functions, ensuring balanced learning across tasks.
  • $\bullet$
    Scalable algorithms with $O(1)$ time and memory cost. We develop Bilevel Loss Balancing for Multi-Task Learning (BiLB4MTL), a highly efficient algorithm tailored to solve the proposed bilevel loss balancing problem. Unlike traditional bilevel methods, BiLB4MTL has a fully single-loop structure without any second-order gradient computation, resulting in an overall $\mathcal{O}(1)$ time and memory complexity. The $2$-task toy example in Figure 1 illustrates that our BiLB4MTL method achieves a more balanced solution compared to other competitive approaches while maintaining superior computational efficiency.
  • $\bullet$
    Superior empirical performance. Extensive experiments demonstrate that our proposed BiLB4MTL method achieves state-of-the-art performance compared to various scalarization-based and gradient manipulation methods across multiple supervised multi-task datasets, including QM9 (Ramakrishnan et al., 2014), CelebA (Liu et al., 2015), and Cityscapes (Cordts et al., 2016). Moreover, BiLB4MTL stands out as one of the most efficient and scalable methods.
  • $\bullet$
    Theoretical guarantees. Theoretically, we show that BiLB4MTL guarantees convergence not only to a stationary point of the bilevel loss balancing problem but also to an $\epsilon$-accurate Pareto stationary point for all $K$ individual loss functions.
Figure 2: Our bilevel loss balancing pipeline for multi-task learning. First, task losses will be normalized through an initial loss normalization module. Then, the lower-level problem optimizes the model parameter $x^{t}$ by minimizing the weighted sum of task losses and the upper-level problem optimizes the router model parameter $W^{t}$ for task balancing.
2 Related Works
Multi-task learning. MTL has recently garnered significant attention in practical applications. One line of research focuses on model architecture, specifically designing various sharing mechanisms (Kokkinos, 2017; Ruder et al., 2019). Another direction addresses the mismatch in loss magnitudes across tasks, proposing methods to balance them. For example, Kendall et al. 2018 balanced tasks by weighting loss functions based on homoscedastic uncertainties, while Liu et al. 2019 dynamically adjusted weights by considering the rate of change in loss values for each task.
Besides, one prominent approach frames MTL as a Multi-Objective Optimization (MOO) problem. Sener & Koltun 2018 introduced this perspective in deep learning, inspiring methods based on the Multi-Gradient Descent Algorithm (MGDA) (Désidéri, 2012). Subsequent work has aimed to address gradient conflicts. For instance, PCGrad (Yu et al., 2020) resolves conflicts by projecting gradients onto the normal plane, GradDrop (Chen et al., 2020) randomly drops conflicting gradients, and CAGrad (Liu et al., 2021a) constrains update directions to balance gradients. Additionally, Nash-MTL (Navon et al., 2022) formulates MTL as a bargaining game among tasks, while FairGrad (Ban & Ji, 2024) incorporates $\alpha$-fairness into gradient adjustments.
On the theoretical side, Zhou et al. 2022 analyzed the convergence properties of stochastic MGDA, and Fernando et al. 2023 proposed a method to reduce bias in the stochastic MGDA with theoretical guarantees. More recent advancements include a double-sampling strategy with provable guarantees introduced by Xiao et al. 2024 and Chen et al. 2024, where the latter one analyzes stochastic MOO algorithms, addressing optimization, generalization, and conflict mitigation trade-offs. Furthermore, Zhang et al. 2024 studied the convergence analysis of both deterministic and stochastic MGDA under a more relaxed generalized smoothness assumption.
Bilevel optimization. Bilevel optimization, first introduced by Bracken & McGill 1973, has been extensively studied over the past few decades. Early research primarily treated it as a constrained optimization problem (Hansen et al., 1992; Shi et al., 2005). More recently, gradient-based methods have gained prominence due to their effectiveness in machine learning applications. Many of these approaches approximate the hypergradient using either linear systems (Domke, 2012; Ji et al., 2021) or automatic differentiation techniques (Maclaurin et al., 2015; Franceschi et al., 2017). However, these methods become impractical in large-scale settings due to their significant computational cost (Xiao & Ji, 2023; Yang et al., 2024b). The primary challenge lies in the high cost of gradient computation: approximating the Hessian-inverse vector requires multiple first- and second-order gradient evaluations, and the nested sub-loops exacerbate this inefficiency. To address these limitations, recent studies have focused on reducing the computational burden of second-order gradients. For example, some methods reformulate the lower-level problem using value-function-based constraints and solve the corresponding Lagrangian formulation (Kwon et al., 2023; Yang et al., 2024a). The work studies convex bilevel problems and proposes a zeroth-order optimization method with finite-time convergence to the Goldstein stationary point (Chen et al., 2023). In this work, we propose a simplified first-order bilevel method for MTL, motivated by intriguing empirical findings.
3 Preliminary
Scalarization-based methods. Multi-task learning (MTL) aims to optimize multiple tasks (objectives) simultaneously with a single model. The straightforward approach is to optimize a weighted summation of all loss functions:
$\min_{x}L_{total}(x)=\sum_{i=1}^{K}w_{i}l_{i}(x),$
where $x\in\mathbb{R}^{d}$ denotes the model parameter, $l_{i}(x):\mathbb{R}^{d}\rightarrow\mathbb{R}_{\geq 0}$ represents the loss function of the $i$-th task and $K$ is the number of tasks. This approach faces two key challenges: 1) fixed weights can lead to significant gradient conflicts, potentially allowing one task to dominate the learning process (Xiao et al., 2024; Wang et al., 2024); and 2) the overall performance is highly sensitive to the weighting of different losses (Kendall et al., 2018). Consequently, such methods often struggle with performance imbalances across tasks.
Gradient manipulation methods. To mitigate gradient conflicts, gradient manipulation methods dynamically compute an update $d^{t}$ at each epoch to balance progress across tasks, where $t$ is the epoch index. The update $d^{t}$ is typically a convex combination of task gradients, expressed as:
$d^{t}=G(x^{t})w^{t},\;\;\text{where}\;w^{t}=h(G(x^{t})),$
with $G(x^{t})=[\nabla l_{1}(x^{t}),\nabla l_{2}(x^{t}),\dots,\nabla l_{K}(x^{t})].$ The weight vector $w^{t}$ is determined by a function $h(\cdot):\mathbb{R}^{K\times d}\rightarrow\mathbb{R}^{K}$, which varies depending on the specific method. However, these methods often require computing and storing the gradients of all $K$ tasks during each epoch, making them less scalable and resource-intensive, particularly in large-scale scenarios. Therefore, it is highly demanding to develop lightweight methods that achieve balanced performance.
Pareto concepts. Solving the MTL problem is challenging because it is difficult to identify a common $x$ that achieves the optima for all tasks. Instead, a widely accepted target is finding a Pareto stationary point. Suppose we have two points $x_{1}$ and $x_{2}$. It is claimed that $x_{1}$ dominates $x_{2}$ if $l_{i}(x_{1})\leq l_{i}(x_{2})\;\forall i\in[K]$, and $\exists j\;l_{j}(x_{1})<l_{j}(x_{2})$. A point is Pareto optimal if it is not dominated by any other points, implying that no task can be improved further without sacrificing another. Besides, a point $x$ is a Pareto stationary point if $\min_{w\in\mathcal{W}}\|G(x)w\|=0$.
4 Bilevel Loss Balancing for Multi-Task Learning
In this section, we present our bilevel loss balancing framework for multi-task learning. As illustrated in Figure 2, this framework contains an initial loss normalization module, a bilevel loss balancing procedure, and a simplified first-order optimization design.
4.1 Initial loss normalization
Before applying loss balancing, an initial loss normalization step is introduced to ensure that the rescaled loss functions are on a similar and comparable scale. This step is necessary because training often involves tasks of different types (e.g., classification and regression) with distinct loss functions, as well as targets measured in varying units or scales (e.g., meters, centimeters, or millimeters). Here, we present three effective approaches that work well in different scenarios.
Identical normalization. This approach maintains the original loss values, where $\tilde{l}_{i}=l_{i},\forall i\in[K]$. It is commonly used in multi-task classification scenarios where the loss functions of all tasks are on a similar scale, such as cross-entropy functions in classification problems.
Rescaled normalization. This method normalizes loss values by rescaling each task’s loss using its initial loss value $l_{i}^{\prime}$, such that $\tilde{l}_{i}=\frac{l_{i}}{l_{i}^{\prime}}$. The resulting normalized loss reflects the training progress and ensures comparability across tasks. This approach is particularly well-suited for scenarios where the loss scales do not differ significantly.
Figure 3: Curves of loss values during the training process for all 11 tasks on the QM9 dataset. The loss values vary significantly across different tasks.
Logarithmic normalization. In some cases, the loss can vary significantly in scale. For example, as shown in Figure 3, we observed that the loss values during training across 9 regression tasks in the QM9 dataset exhibit substantial differences in magnitude, with loss ratios exceeding 1000 in certain instances. To address this issue, we propose logarithmically rescaling the loss functions such that $\tilde{l}_{i}=\log\left(\frac{l_{i}}{l_{i,0}}\right)$, where $l_{i,0}$ represents the initial loss value for the $i$-th task at each epoch. Our experiments demonstrate that this initialization approach stabilizes training by reducing large fluctuations caused by significant scale variations.
4.2 Bilevel loss balancing formulation
Building on the normalized loss function values, we introduce a novel bilevel loss balancing approach for MTL to achieve a more balanced and fair solution. The formulation is as follows:
$\min_{W}\sum_{i=1}^{K-1}|\tau_{i}\tilde{l}_{i}(x^{*})-\tau_{i+1}\tilde{l}_{i+1}(x^{*})|:=f(W,x^{*})$
[…]
where we denote $x^{*}=x^{*}(W)$ for notational convenience. It can be seen from section 4.2 that The lower-level problem minimizes the weighted sum of losses w.r.t. the model parameters $x$, while the upper-level problem minimizes the accumulated weighted loss gaps w.r.t. the parameters $W$, ensuring balance across different tasks. Note that we define a routing function $\sigma(W)\in\mathbb{R}^{K}$, which is parameterized by a small neural network with a softmax output layer. For the upper-level weight vector $\tau=(\tau_{1},...,\tau_{K})$, we provide two effective options: (i) $\tau=\sigma(W)$ and (ii) $\tau={\mathbf{1}}$ that work well in the experiments.
4.3 Scalable first-order algorithm design
To enable large-scale applications, we adopt an efficient and scalable first-order method to solve the problem in Section 4.2. Inspired by recent advancements in first-order bilevel optimization (Kwon et al., 2023; Yang et al., 2023), we reformulate the original bilevel problem into an equivalent constrained optimization problem as follows.
$\min_{W}f(W,x)\;\;\text{s.t. }\underbrace{\sum_{i=1}^{K}\sigma_{i}(W)\tilde{l}_{i}(x)-\sum_{i=1}^{K}\sigma_{i}(W)\tilde{l}_{i}(x^{*})}_{\text{penalty function }p(W,x)}\leq 0.$
Then, given a penalty constant $\lambda>0$, penalizing $p(W,x)$ into the upper-level loss function yields
$\min_{W,x}f(W,x)+\lambda\sum_{i=1}^{K}(\sigma_{i}(W)\tilde{l}_{i}(x)-\sigma_{i}(W)\tilde{l}_{i}(x^{*})).$ (2)
Intuitively, a larger $\lambda$ allows more precise training on model parameters $x$ such that $x$ converges closer to $x^{*}$. Conversely, a smaller $\lambda$ prioritizes upper-level loss balancing during training. The main challenge of solving the penalized problem above lies in the updates of $W$, as shown below:
$W^{t+1}=$ $W^{t}-\alpha(\nabla_{W}f(W^{t},x^{t})+$
$\lambda(\nabla_{W}g(W^{t},x^{t})-\nabla_{W}g(W^{t},z_{N}^{t}))),$ (3)
where $t$ is the epoch index, $\alpha$ is the step size, and $z_{N}^{t}$ is an approximation of $x_{t}^{*}\in\arg\min_{x}g(W^{t},x)$ through the following loop of $N$ iterations each epoch.
$z_{n+1}^{t}=z_{n}^{t}-\beta\nabla_{z}g(W^{t},z_{n}^{t}),n=0,1,...N-1,$ (4)
where $N$ is typically chosen to be sufficiently large, ensuring that $z_{N}^{t}$ closely approximates $x_{t}^{*}$ (full algorithm is provided in Algorithm 2 in the appendix). Consequently, this sub-loop of iterations incurs significant computational overhead, driven by the high dimensionality of $z$ (matching that of the model parameters) and the large value of $N$.
Interestingly, our experiments reveal that the gradient norm $\|\nabla_{W}g(W^{t},z_{N}^{t})\|$ remains sufficiently small, typically orders of magnitude smaller than the gradient norm $\|\nabla_{W}g(W^{t},x^{t})\|$, which is used to update outer parameters $W$. This behavior is illustrated in Figure 4. Specifically, we set $N=50$ during training. On average, the ratio $\|\nabla_{W}g(W^{t},x^{t})\|/\|\nabla_{W}g(W^{t},z_{N}^{t})\|$ exceeds 100, despite some fluctuations. Under these conditions, the term $\nabla_{W}g(W,z_{N}^{t})$ can be safely neglected, thereby eliminating the need for the expensive loop
Figure 4: Gradient norm values during the training process on the Cityscapes dataset. Similar phenomena have also been observed in other datasets.
in Equation 4. This approximation has been effectively utilized in large-scale applications, such as fine-tuning large language models, to reduce memory and computational costs (Shen et al., 2024a). It also serves as a foundation for our proposed algorithm, Bilevel Loss Balancing for Multi-Task Learning (BiLB4MTL), described in Algorithm 1. BiLB4MTL employs a fully single-loop structure, requiring only a single gradient computation for both variables per epoch, resulting in an $\mathcal{O}(1)$ time and memory cost.
Algorithm 1 BiLB4MTL
1:  Initialize: $W^{0},x^{0}$
2:  for $t=0,1,...,T-1$ do
3:     $x^{t+1}=x^{t}-\alpha(\nabla_{x}f(W^{t},x^{t})+\lambda\nabla_{x}g(W^{t},x^{t}))$.
4:     $W^{t+1}=W^{t}-\alpha(\nabla_{W}f(W^{t},x^{t})+\lambda\nabla_{W}g(W^{t},x^{t}))$
5:  end for
In Section 6, we show that our BiLB4MTL method attains both an $\epsilon$-accurate stationary point for the bilevel problem in section 4.2 and an $\epsilon$-accurate Pareto stationary point for original loss functions under mild conditions.
5 Empirical Results
In this section, we conduct extensive practical experiments under multi-task classification, regression, and mixed settings to demonstrate the effectiveness of our method. Full experimental details can be found in Appendix A.
Table 1: Results on Cityscapes (2-task) dataset. Each experiment is repeated 3 times with different random seeds and the average is reported. The best results are highlighted in bold, while the second-best results are indicated with underlines.
Method Segmentation Depth $\Delta m\%\downarrow$ mIoU $\uparrow$ Pix Acc $\uparrow$ Abs Err $\downarrow$ Rel Err $\downarrow$ STL 74.01 93.16 0.0125 27.77 LS 75.18 93.49 0.0155 46.77 22.60 SI 70.95 91.73 0.0161 33.83 14.11 RLW 74.57 93.41 0.0158 47.79 24.38 DWA 75.24 93.52 0.0160 44.37 21.45 UW 72.02 92.85 0.0140 30.13 5.89 FAMO 74.54 93.29 0.0145 32.59 8.13 GO4Align 72.63 93.03 0.0164 27.58 8.11 MGDA 68.84 91.54 0.0309 33.50 44.14 PCGrad 75.13 93.48 0.0154 42.07 18.29 GradDrop 75.27 93.53 0.0157 47.54 23.73 CAGrad 75.16 93.48 0.0141 37.60 11.64 IMTL-G 75.33 93.49 0.0135 38.41 11.10 MoCo 75.42 93.55 0.0149 34.19 9.90 Nash-MTL 75.41 93.66 0.0129 35.02 6.82 FairGrad 75.72 93.68 0.0134 32.25 5.18 BiLB4MTL ($\tau=\bf 1$) 74.53 93.42 0.0128 26.79 -0.57 BiLB4MTL ($\tau=\sigma$) 74.92 93.24 0.0127 29.80 1.93
Baselines and evaluation. To demonstrate the effectiveness of our proposed method, we evaluate its performance against a broad range of baseline approaches. The compared methods include scalarization-based algorithms, such as Linear Scalarization (LS), Scale-Invariant (SI), Random Loss Weighting (RLW) (Lin et al., 2021), Dynamic Weight Average (DWA) (Liu et al., 2019), and Uncertainty Weighting (UW) (Kendall et al., 2018), GO4Align (Shen et al., 2024b). We also benchmark against gradient manipulation methods, including Multi-Gradient Descent Algorithm (MGDA) (Désidéri, 2012), PCGrad (Yu et al., 2020), GradDrop (Chen et al., 2020), CAGrad (Liu et al., 2021a), IMTL-G (Liu et al., 2021b), Nash-MTL (Navon et al., 2022), FAMO (Liu et al., 2024), and FairGrad (Ban & Ji, 2024). To provide a comprehensive evaluation, we report the performance of each individual task and employ one additional metric: $\bf\Delta m\%$ to quantify overall performance. The $\bf\Delta m\%$ metric measures the average relative performance drop of a multi-task model compared to its corresponding single-task baseline. Formally, it is defined as:
$\Delta m\%=\frac{1}{K}\sum_{i=1}^{K}(-1)^{\delta_{k}}(M_{m,k}-M_{b,k})/M_{b,k}\times 100,$
where $M_{m,k}$ and $M_{b,k}$ represent the performance of the $k$-th task for the multi-task model $m$ and single-task model $b$, respectively. The indicator $\delta_{k}=1$ if lower values indicate better performance and 0 otherwise.
5.1 Experiment setup
Here, we summarize our experimental setup, including the dataset, configuration, and hyperparameter tuning.
Image-Level Classification. CelebA (Liu et al., 2015), one of the most widely used datasets, is a large-scale facial attribute dataset containing over 200K celebrity images. Each image is annotated with 40 attributes, such as the presence of eyeglasses and smiling. Following the experimental setup in Ban & Ji 2024, we treat CelebA as a 40-task multi-task learning (MTL) classification problem, where each task predicts the presence of a specific attribute. Since all tasks involve binary classification with the same binary cross-entropy loss function, we apply identical normalization for both options of $\tau=\bf 1$ and $\tau=\sigma$ at the initial loss normalization stage. The network architecture consists of a 9-layer convolutional neural network (CNN) as the shared model, with multiple linear layers serving as task-specific heads. We train the model for 15 epochs using the Adam optimizer with a batch size of 256.
Regression. QM9 (Ramakrishnan et al., 2014) dataset is another widely used benchmark for multi-task regression problems in quantum chemistry. It contains 130K molecules represented as graphs, and 11 properties to be predicted. Though all tasks share the same loss function, mean squared error, they exhibit varying scales: a phenomenon commonly observed in regression tasks but less prevalent in classification tasks, as shown in Figure 3. To address this scale discrepancy, we adopt the logarithmic normalization in Section 4.1 at the initial loss normalization stage for both options of $\tau=\bf 1$ and $\tau=\sigma$. Following the experimental setup in Liu et al. 2024; Navon et al. 2022, we use the same model and data split, 110K molecules for training, 10k for validation, and the rest 10k for testing. The model is trained 300 epochs with a batch size of 120. The learning rate starts at 1e-3 and is reduced whenever the validation performance stagnates for 5 consecutive epochs.
Table 2: Results on CelebA (40-task) and NYU-v2 (3-task) datasets. Each experiment is repeated 3 times and the average is reported. The best results are highlighted in bold, while the second-best results are indicated with underlines.
Method CelebA (40 tasks) NYU-v2 (3 tasks) $\Delta m\%\downarrow$ $\Delta m\%\downarrow$ LS 4.15 5.59 SI 7.20 4.39 RLW 1.46 7.78 DWA 3.20 3.57 UW 3.23 4.05 FAMO 1.21 -4.10 GO4Align 0.88 -6.08 MGDA 14.85 1.38 PCGrad 3.17 3.97 CAGrad 2.48 0.20 IMTL-G 0.84 -0.76 Nash-MTL 2.84 -4.04 FairGrad 0.37 -4.66 BiLB4MTL ($\tau=\bf 1$) -1.07 -4.40 BiLB4MTL ($\tau=\sigma$) -1.31 -3.52
Dense Prediction. The Cityscapes dataset (Cordts et al., 2016) consists of 5000 street-scene images designed for two tasks: 7-class semantic segmentation (a classification task) and depth estimation (a regression task). Similarly, the NYU-v2 dataset (Silberman et al., 2012) is widely used for indoor scene understanding and contains 1449 densely annotated images. It includes one pixel-level classification task, semantic segmentation, and two pixel-level regression tasks, 13-class depth estimation plus surface normal prediction. These datasets provide benchmarks for evaluating the performance of our method in mixed multi-task settings. Since the number of tasks is small and the loss values exhibit minimal variation, we applied rescaled normalization when selecting $\tau=\sigma$ and identical normalization when selecting $\tau=\bf 1$. We follow the same experimental setup described in Liu et al. 2021a; Navon et al. 2022 and adopt MTAN (Liu et al., 2019) as the backbone, which incorporates task-specific attention modules into SegNet (Badrinarayanan et al., 2017). Both models are trained for 200 epochs, with batch sizes of 8 for Cityscapes and 8 for NYU-v2. The learning rates are initialized at 3e-4 and 1e-4 for the first 100 epochs and reduced by half for the remaining epochs, respectively.
Figure 5: Time scale comparison among well-performing approaches, with LS considered the reference method for standard time.
Table 3: Detailed results of on QM9 (11-task) dataset. Each experiment is repeated 3 times and the average is reported. The best results are highlighted in bold, while the second-best results are indicated with underlines.
Method $\mu$ $\alpha$ $\epsilon_{HOMO}$ $\epsilon_{LUMO}$ $\langle R^{2}\rangle$ ZPVE $U_{0}$ $U$ $H$ $G$ $c_{v}$ $\Delta m\%\downarrow$ MAE $\downarrow$ STL 0.067 0.181 60.57 53.91 0.502 4.53 58.8 64.2 63.8 66.2 0.072 LS 0.106 0.325 73.57 89.67 5.19 14.06 143.4 144.2 144.6 140.3 0.128 177.6 SI 0.309 0.345 149.8 135.7 1.00 4.50 55.3 55.75 55.82 55.27 0.112 77.8 RLW 0.113 0.340 76.95 92.76 5.86 15.46 156.3 157.1 157.6 153.0 0.137 203.8 DWA 0.107 0.325 74.06 90.61 5.09 13.99 142.3 143.0 143.4 139.3 0.125 175.3 UW 0.386 0.425 166.2 155.8 1.06 4.99 66.4 66.78 66.80 66.24 0.122 108.0 FAMO 0.15 0.30 94.0 95.2 1.63 4.95 70.82 71.2 71.2 70.3 0.10 58.5 GO4Align 0.17 0.35 102.4 119.0 1.22 4.94 53.9 54.3 54.3 53.9 0.11 52.7 MGDA 0.217 0.368 126.8 104.6 3.22 5.69 88.37 89.4 89.32 88.01 0.120 120.5 PCGrad 0.106 0.293 75.85 88.33 3.94 9.15 116.36 116.8 117.2 114.5 0.110 125.7 CAGrad 0.118 0.321 83.51 94.81 3.21 6.93 113.99 114.3 114.5 112.3 0.116 112.8 IMTL-G 0.136 0.287 98.31 93.96 1.75 5.69 101.4 102.4 102.0 100.1 0.096 77.2 Nash-MTL 0.102 0.248 82.95 81.89 2.42 5.38 74.5 75.02 75.10 74.16 0.093 62.0 FairGrad 0.117 0.253 87.57 84.00 2.15 5.07 70.89 71.17 71.21 70.88 0.095 57.9 BiLB4MTL ($\tau=\bf 1$) 0.341 0.405 161.49 140.06 1.65 5.04 75.31 74.82 75.66 75.82 0.125 113.6 BiLB4MTL ($\tau=\sigma$) 0.23 0.29 123.89 111.95 0.97 3.99 42.73 43.1 43.2 43.1 0.097 49.5
Hyperparameter tuning. In our method, hyperparameters include the step size $\alpha$ and the penalty constant $\lambda$. For the step size, we adopt the settings from prior experiments without extensive tuning. While we use the same step size for updates to both $W$ and $x$ in our implementation, these can be adjusted independently in practice. For $\lambda$, we determine the optimal value through a grid search. Starting with a coarse range $\lambda\in[0.01,0.1,1,2,5,10]$, we evaluate performance and determine which value is better. Then we narrow down the search space and continue to execute a finer grid search with a step size of $0.1$ or $0.02$ until we determine an appropriate value.
5.2 Experimental results
Results on the four benchmark datasets are provided in Table 1, Table 2, and Table 3. We observe that BiLB4MTL outperforms existing methods on both the CelebA and QM9 datasets, achieving the lowest performance drops of $\Delta m\%=-1.31$ and $\Delta m\%=49.5$, respectively. Detailed results for the QM9 dataset illustrate that it achieves a balanced performance across all tasks. Besides, the BiLB4MTL($\tau=\bf 1$) does not effectively minimize task disparities due to large loss scale differences. These results highlight the effectiveness of our method in handling a large number of tasks in both classification and regression settings. Meanwhile, it achieves the lowest performance drop, with $\Delta m\%=-0.57$ on the Cityscapes dataset, while delivering comparable results on the NYU-v2 dataset, where the detailed results are shown in Table 4 in the Appendix. These findings highlight the capability of BiLB4MTL to effectively handle mixed multi-task learning scenarios.
Additionally, we conducted an ablation study to evaluate the impact of different normalization strategies. The results, presented in Table 5 in the Appendix, demonstrate the effectiveness of our proposed normalization methods.
5.3 MTL efficiency
Finally, we compare the running times of well-performing approaches in Figure 5. Notably, our BiLB4MTL introduces negligible overhead compared with LS with at most a $1\times1$ increase, aligning with other $\mathcal{O}(1)$ methods such as GO4Align and FAMO. In contrast, gradient manipulation methods, taking $\mathcal{O}(K)$ computational cost, become significantly slower in many-task scenarios. For example, Nash-MTL requires approximately 12×12$\times$12 × times more than BiLB4MTL on the CelebA dataset.
6 Theoretical Analysis
In this section, we provide convergence analysis for our BiLB4MTL method. We first provide several useful definitions and assumptions.
Definition 1.
Given $L>0$, a function $\ell$ is said to be $L$-Lipschitz-continuous on $\mathcal{X}$ if it holds for any $x,x^{\prime}\in\mathcal{X}$ that $\|\ell(x)-\ell(x^{\prime})\|\leq L\|x-x^{\prime}\|$. A function $\ell$ is said to be $L$-Lipschitz-smooth if its gradient is $L$-Lipschitz-continuous.
Definition 2 (Pareto stationarity).
We say $x$ is an $\epsilon$-accurate Pareto stationary point for loss functions $\{l_{i}(x)\}$ if $\min_{w\in\mathcal{W}}\|G(x)w\|^{2}=\mathcal{O}(\epsilon)$, where $G(x)=[\nabla l_{1}(x),\nabla l_{2}(x),...,\nabla l_{K}(x)]$.
Inspired by Shen & Chen 2023, we define the following two surrogates of the original bilevel problem in section 4.2, as shown below.
Definition 3.
Define two surrogate bilevel problems as
$\mathcal{BP}_{\lambda}:$ $\min_{W,x}f(W,x)+\lambda(g(W,x)-g(W,x^{*})),$
$\mathcal{BP}_{\epsilon}:$ $\min_{W,x}f(W,x)\;\;\text{s.t. }g(W,x)-g(W,x^{*})\leq\epsilon,$ (5)
where $\mathcal{BP}_{\lambda}$ is the penalized bilevel problem, and $\mathcal{BP}_{\epsilon}$ recovers to the original bilevel problem if $\epsilon=0$.
Assumption 1 (Lipschitz and smoothness).
There exists a constant $L$ such that the upper-level function $f(W,\cdot)$ is $L$-Lipschitz continuous. There exists constants $L_{f}$ and $L_{g}$ such that functions $f(W,x)$ and $g(W,x)$ are $L_{f}$- and $L_{g}$-Lipschitz-smooth.
Assumption 2 (Polyak-Lojasiewicz (PL) condition).
The lower-level function $g(W,\cdot)$ satisfies the $\frac{1}{\mu}$-PL condition such that given any $W$, the following inequality holds for any feasible $x$.
$\|\nabla g(W,x)\|^{2}\geq\frac{1}{\mu}(g(W,x)-g(W,x^{*})).$
Lipschitz continuity and smoothness are standard assumptions in the study of bilevel optimization (Ghadimi & Wang, 2018; Ji et al., 2021). While the absolute values in the upper-level function in Section 4.2 are non-smooth, they can be easily modified to ensure smoothness, such as by using a soft absolute value function of the form $y=\sqrt{x^{2}+\epsilon}$ where $\epsilon$ is a small positive constant. Moreover, the PL condition can be satisfied in over-parameterized neural network settings (Mei et al., 2020; Frei & Gu, 2021). The following theorem presents the convergence analysis of our algorithms.
Theorem 1.
Suppose 1-2 are satisfied. Select hyperparameters
$\alpha\in\Big{(}0,\frac{1}{L_{f}+\lambda(2L_{g}+L_{g}^{2}\mu)}\Big{]},\;\beta\in(0,\frac{1}{L_{g}}],$
$\lambda=L\sqrt{3\mu{\epsilon}^{-1}},\text{ and }N=\Omega(\log(\alpha t)).$
(i) Our method with the updates section 4.3 and eq. 4 (i.e., Algorithm 2 in the appendix) finds an $\epsilon$-accurate stationary point of the problem $\mathcal{BP}_{\lambda}$. If this stationary point is a local/global solution to $\mathcal{BP}_{\lambda}$, it is also a local/global solution to $\mathcal{BP}_{\epsilon}$. Furthermore, it is also an $\epsilon$-accurate Pareto stationary point for loss functions $l_{i}(x),i=1,...,K.$
(ii) Moreover, if $\|\nabla_{W}g(W^{t},z_{N}^{t})\|=\mathcal{O}(\epsilon)$ for $t=1,...,T$. The simplified method in Algorithm 1 also achieves the same convergence guarantee as that in $(i)$ with $\mathcal{O}(\epsilon^{-2})$ iterations.
The complete proof is provided in Theorem 2. In the first part of our theorem, we establish a connection between the stationarity of $\mathcal{BP}_{\lambda}$ and Pareto stationarity, as well as an equivalence between the Pareto stationarities of the original loss functions $\{l_{i}\}$ and the normalized loss functions $\{\tilde{l}_{i}\}$. The second part of Theorem 1 introduces an additional gradient vanishing assumption, which has been validated in our experiments. It demonstrates that our simplified BiLB4MTL method can also attain an $\epsilon$-accurate stationary point for the problem $\mathcal{BP}_{\lambda}$ and an $\epsilon$-accurate Pareto stationary point for the original loss functions.
7 Conclusion
We introduced BiLB4MTL, a scalable loss balancing approach for multi-task learning based on bilevel optimization. Our method achieves efficient loss balancing with only $\mathcal{O}(1)$ time and memory complexity while guaranteeing convergence to both a stationary point of the bilevel problem and an $\epsilon$-accurate Pareto stationary point for all task loss functions. Extensive experiments demonstrate that BiLB4MTL outperforms existing methods in both accuracy and efficiency, highlighting its effectiveness for large-scale MTL.
For future work, we plan to explore the application of our method to broader multi-task learning problems, including recommendation systems.
Impact Statement
This paper presents work to advance the field of multi-task Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Experiment setup
A.1 Toy example
To better understand the benefits of our method, we illustrate the training trajectory along with the training time in a toy example of 2-task learning following the same setting in FAMO (Liu et al., 2024). The loss functions $L_{1}(x),L_{2}(x)$, where $x$ is the model parameter, of two tasks are listed below.
$L_{1}(x)$ $=0.1\times(c_{1}(x)f_{1}(x)+c_{2}(x)g_{1}(x)),\;\;L_{2}(x)=c_{1}(x)f_{2}(x)+c_{2}(x)g_{2}(x)\;\;\text{where}$
$f_{1}(x)$ $=\log{\big{(}\max(|0.5(-x_{1}-7)-\tanh{(-x_{2})}|,~{}~{}0.000005)\big{)}}+6,$
$f_{2}(x)$ $=\log{\big{(}\max(|0.5(-x_{1}+3)-\tanh{(-x_{2})}+2|,~{}~{}0.000005)\big{)}}+6,$
$g_{1}(x)$ $=\big{(}(-x_{1}+7)^{2}+0.1*(-x_{2}-8)^{2}\big{)}/10-20,$
$g_{2}(x)$ $=\big{(}(-x_{1}-7)^{2}+0.1*(-x_{2}-8)^{2})\big{/}10-20,$
$c_{1}(x)$ $=\max(\tanh{(0.5*x_{2})},~{}0)~{}~{}\text{and}~{}~{}c_{2}(x)=\max(\tanh{(-0.5*x_{2})},~{}0).$ (6)
In Figure 1, the black dots represent 5 chosen initial points $\{(-8.5,7.5),(-8.5,5),(0,0),(9,9),(10,-8)\}$ while the black stars represent the converging points on the Pareto front. We use the Adam optimizer and train each method for 50k steps. Our method can always converge to balanced results efficiently. We use Adam optimizer with a learning rate of 1e-3. The training time is recalculated according to real-time ratios in our machine. We find that LS and MGDA do not converge to balanced points while FAMO converges to balanced results to some extent. Meanwhile, our method with rescale normalization can always converge to balanced results efficiently.
Table 4: Results on NYU-v2 (3-task) dataset. Each experiment is repeated 3 times with different random seeds and the average is reported.
Method Segmentation Depth Surface Normal $\Delta m\%\downarrow$ mIoU $\uparrow$ Pix Acc $\uparrow$ Abs Err $\downarrow$ Rel Err $\downarrow$ Angle Distance $\downarrow$ Within $t^{\circ}$uparrow Mean Median 11.25 22.5 30 STL 38.30 63.76 0.6754 0.2780 25.01 19.21 30.14 57.20 69.15 LS 39.29 65.33 0.5493 0.2263 28.15 23.96 22.09 47.50 61.08 5.59 SI 38.45 64.27 0.5354 0.2201 27.60 23.37 22.53 48.57 62.32 4.39 RLW 37.17 63.77 0.5759 0.2410 28.27 24.18 22.26 47.05 60.62 7.78 DWA 39.11 65.31 0.5510 0.2285 27.61 23.18 24.17 50.18 62.39 3.57 UW 36.87 63.17 0.5446 0.2260 27.04 22.61 23.54 49.05 63.65 4.05 FAMO 38.88 64.90 0.5474 0.2194 25.06 19.57 29.21 56.61 68.98 -4.10 GO4Align 40.42 65.37 0.5492 0.2167 24.76 18.94 30.54 57.87 69.84 -6.08 MGDA 30.47 59.90 0.6070 0.2555 24.88 19.45 29.18 56.88 69.36 1.38 PCGrad 38.06 64.64 0.5550 0.2325 27.41 22.80 23.86 49.83 63.14 3.97 GradDrop 39.39 65.12 0.5455 0.2279 27.48 22.96 23.38 49.44 62.87 3.58 CAGrad 39.79 65.49 0.5486 0.2250 26.31 21.58 25.61 52.36 65.58 0.20 IMTL-G 39.35 65.60 0.5426 0.2256 26.02 21.19 26.20 53.13 66.24 -0.76 MoCo 40.30 66.07 0.5575 0.2135 26.67 21.83 25.61 51.78 64.85 0.16 Nash-MTL 40.13 65.93 0.5261 0.2171 25.26 20.08 28.40 55.47 68.15 -4.04 FairGrad 39.74 66.01 0.5377 0.2236 24.84 19.60 29.26 56.58 69.16 -4.66 BiLB4MTL($\tau=\bf 1$) 38.04 38.04 0.5402 0.2278 24.70 19.19 29.97 57.44 69.69 -4.40 BiLB4MTL($\tau=\sigma$) 40.94 66.21 0.5316 0.2210 25.33 20.35 27.79 54.95 67.82 -3.52
A.2 Ablation study on normalization alternatives
We conducted an ablation study on normalization strategies for the QM9 dataset, where the $\Delta m\%$ metric was computed and presented in Table 5. The best results were achieved by using a weighted sum of log-normalized losses as the lower-level objective and a weighted sum of log-normalized loss gaps as the upper-level objective. This approach aligns with our formulation in Section 4.2, incorporating logarithmic normalization of loss values. In Table 5, weighted indicates the inclusion of $\sigma(W)$, log refers to the application of the logarithmic operator, and normalized signifies the division by the initial loss values.
Table 5: Ablation study on normalization alternatives.
Upper-Level Lower-Level Weighted normalized loss Weighted log-normalized loss
Weighted log-normalized loss gaps 150.3 49.5
Log-normalized loss gaps 103.9 113.6
Weighted normalized loss gaps 197.3 94.5
Normalized loss gaps 147.5 157.6
Weighted Loss gaps 125.8 72.6
Loss gaps 172.3 76.7
Appendix B Additional information
Here, we present the complete version of the double-loop algorithm for solving the penalized bilevel problem, $\mathcal{BP}_{\lambda}$ in Equation 2, as detailed in Algorithm 2. Notably, the local or global solution of $\mathcal{BP}_{\lambda}$ obtained by Algorithm 2 also serves as a local or global solution to $\mathcal{BP}_{\epsilon}$, as established by Proposition 2 in Shen & Chen 2023.
Algorithm 2 Double-loop First-order method
1:  Initialize: $W^{0},x^{0},z_{0}^{0}$
2:  for $t=0,1,...,T-1$ do
3:     warm start $z_{0}^{t}=x^{t}$
4:     for $n=0,1,...,N$ do
5:        $z_{n+1}^{t}=z_{n}^{t}-\beta\lambda\nabla_{z}g(W^{t},z_{n}^{t})$
6:     end for
7:     $x^{t+1}=x^{t}-\alpha(\nabla_{x}f(W^{t},x^{t})+\lambda\nabla_{x}g(W^{t},x^{t}))$
8:     $W^{t+1}=W^{t}-\alpha(\nabla_{W}f(W^{t},x^{t})+\lambda(\nabla_{W}g(W^{t},x^{t})-\nabla_{W}g(W^{t},z_{N}^{t})))$
9:  end for
Appendix C Analysis
In the analysis, we need the following definitions.
$x_{t}^{*}=$ $\arg\min_{x}g(W^{t},x),\;G(x)=[\nabla l_{1}(x),\nabla l_{2}(x),...,\nabla l_{K}(x)],\widetilde{G}(x)=[\nabla\tilde{l}_{1}(x),\nabla\tilde{l}_{2}(x),...,\nabla\tilde{l}_{K}(x)]$
$F(\theta^{t})=$ $f(\theta^{t})+\lambda p(\theta^{t}),\Phi(\theta^{t})=f(\theta^{t})+\lambda g(\theta^{t}),\;\text{where}\;\theta^{t}=(W^{t},x^{t}),p(\theta^{t})=g(W^{t},x^{t})-g(W^{t},x_{t}^{*})$
$\nabla f(W,x)=$ $(\nabla_{W}f(W,x),\nabla_{x}f(W,x)),\nabla g(W,x)=(\nabla_{W}g(W,x),\nabla_{x}g(W,x)).$ (7)
Lemma 1.
Let $(W,x)$ be a solution to the $\mathcal{BP}_{\epsilon}$. This point is also an $\epsilon$-accurate Pareto stationarity point for $\{l_{i}(x)\}$ satisfying
$\min_{w\in\mathcal{W}}\|G(x)w\|^{2}=\mathcal{O}(\epsilon).$
Proof.
According to the definition of $\mathcal{BP}_{\epsilon}$, its solution $(W,x)$ satisfies that
$g(W,x)-g(W,x^{*})\leq\epsilon.$ (8)
Further, according to 1, we can obtain
$g(W,x)\geq g(W,x^{*})+\nabla_{x}g(W,x^{*})(x-x^{*})+\frac{1}{2L_{g}}\|\nabla_{x}g(W,x)-\nabla_{x}g(W,x^{*})\|^{2}.$
Since $x^{*}\in\arg\min_{x}g(W,x)$ and $g(W,x)=\sum_{i=1}^{K}\sigma_{i}(W)\tilde{l}_{i}(x)$, we have $\nabla_{x}g(W,x^{*})=0$ and $\nabla_{x}g(W,x)=\sum_{i=1}^{K}\sigma_{i}(W)\nabla_{x}\tilde{l}_{i}(x)=\widetilde{G}(x)\sigma(W)$ We can obtain,
$\|\widetilde{G}(x)\sigma(W)\|^{2}\leq 2L_{g}(g(W,x)-g(W,x^{*}))=\mathcal{O}(\epsilon),$ (9)
where the last inequality follows from Equation 8. Furthermore, since we have used softmax at the last layer of our neural network, $\sigma(W)$ belongs to the probability simplex $\mathcal{W}$. Thus we can derive
$\min_{w\in\mathcal{W}}\|\widetilde{G}(x)w\|^{2}\leq\|\widetilde{G}(x)\sigma(W)\|^{2}=\mathcal{O}(\epsilon).$
Thus, the solution $(W,x)$ to the $\mathcal{BP}_{\epsilon}$ also satisfies Pareto stationarity of the normalized loss functions $\{\tilde{l}_{i}\}$. Then we show the equivalence between the Pareto stationarities for the original loss functions $\{l_{i}\}$ and the normalized loss functions $\{\tilde{l}_{i}\}$. First, for the identical normalization, the above eq. 9 naturally recovers
$\min_{w\in\mathcal{W}}\|{G}(x)w\|^{2}=\min_{w\in\mathcal{W}}\|\widetilde{G}(x)w\|^{2}\leq\|\widetilde{G}(x)\sigma(W)\|^{2}=\mathcal{O}(\epsilon).$
Then, for the rescaled normalization, $\tilde{l}_{i}(x)=\frac{l_{i}(x)}{l_{i}^{\prime}}$ where $\forall i,l_{i}^{\prime}=\mathcal{O}(1)$. Thus, we can have
$\min_{w\in\mathcal{W}}\|G(x)w\|^{2}\leq\|G(x)\sigma(W)\|^{2}\leq L^{2}_{\text{max}}\|\widetilde{G}(x)\sigma(W)\|^{2}=\mathcal{O}(\epsilon),$
where $L_{\text{max}}=\max_{i}l_{i}^{\prime}$. Finally, for the logarithmic normalization, $\tilde{l}_{i}(x)=\log\Big{(}\frac{l_{i}(x)}{l_{i,0}}\Big{)}=\lim_{\kappa\rightarrow 1}\frac{\Big{(}\frac{l_{i}(x)}{l_{i,0}}\Big{)}^{1-\kappa}-1}{1-\kappa}$. Furthermore, according to the Proposition 6.1 in Ban & Ji 2024, the Pareto front of the $(\tilde{l}_{1}(x),\tilde{l}_{2}(x),...,\tilde{l}_{K}(x))$ which can be considered as $\kappa$-fair functions is the same as that of loss functions $\Big{(}\frac{l_{1}(x)}{l_{1,0}},\frac{l_{2}(x)}{l_{2,0}},...,\frac{l_{K}(x)}{l_{K,0}}\Big{)}$. Therefore, for an $\epsilon$-accurate Pareto stationarity point $x$ of the normalized loss functions, we can obtain
$\min_{w\in\mathcal{W}}\|G^{\prime}(x)w\|^{2}=\mathcal{O}(\epsilon),$
where $G^{\prime}(x)=\Big{(}\frac{\nabla l_{1}(x)}{l_{1,0}},\frac{\nabla l_{2}(x)}{l_{2,0}},...,\frac{\nabla l_{K}(x)}{l_{K,0}}\Big{)}$. Furthermore, we can obtain
$\min_{w\in\mathcal{W}}\|G(x)w\|^{2}\leq\min_{w\in\mathcal{W}}(L^{\prime}_{\text{max}})^{2}\|G^{\prime}(x)w\|^{2}=\mathcal{O}(\epsilon),$
where $L^{\prime}_{\text{max}}=\max_{i}l_{i,0}=\mathcal{O}(1)$. Then with our three normalization approaches, there is an equivalence between the Pareto stationarities of the original loss functions $\{l_{i}\}$ and the normalized loss functions $\{\tilde{l}_{i}\}$. The proof is complete. ∎
Theorem 2 (Restatement of Theorem 1).
Suppose 1-2 are satisfied. Select hyperparameters
$\alpha\in(0,\frac{1}{L_{f}+\lambda(2L_{g}+L_{g}^{2}\mu)}],\;\beta\in(0,\frac{1}{L_{g}}],\lambda=L\sqrt{3\mu{\epsilon}^{-1}},\text{ and }N=\Omega(\log(\alpha t)).$
(i) Our method with the updates section 4.3 and eq. 4 (i.e., Algorithm 2 in the appendix) finds an $\epsilon$-accurate stationary point of the problem $\mathcal{BP}_{\lambda}$. If this stationary point is a local/global solution to $\mathcal{BP}_{\lambda}$, it is also a local/global solution to $\mathcal{BP}_{\epsilon}$. Furthermore, it is also an $\epsilon$-accurate Pareto stationary point for loss functions $l_{i}(x),i=1,...,K.$
(ii) Moreover, if $\|\nabla_{W}g(W^{t},z_{N}^{t})\|=\mathcal{O}(\epsilon)$ for $t=1,...,T$. The simplified method in Algorithm 1 also achieves the same convergence guarantee as that in $(i)$ with $\mathcal{O}(\epsilon^{-2})$ iterations.
Proof.
We start with the first half of our theorem. Directly from Theorem 3 in Shen & Chen 2023, Algorithm 2 achieves an $\epsilon$-accurate stationary point of $\mathcal{BP}_{\lambda}$ with $\widetilde{\mathcal{O}}(\epsilon^{-1.5})$ iterations such that
$\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla f(W^{t},x^{t})+\lambda(\nabla g(W^{t},x^{t})-\nabla g(W^{t},x_{t}^{*}))\|^{2}\leq\frac{F(W^{0},x^{0})}{\alpha T}+\frac{10L^{2}L_{g}^{2}}{T}=\mathcal{O}(\epsilon).$
Recall that $F(W^{0},x^{0})=f(W^{0},x^{0})+\lambda(g(W^{0},x^{0})-g(W^{0},x_{0}^{*}))$. According to the Proposition 2 in Shen & Chen 2023 by setting $\delta=\epsilon$ therein, we can have $g(W^{T},x^{T})-g(W^{T},x_{T}^{*})\leq\epsilon$ if this stationary point is local/global solution to $\mathcal{BP}_{\lambda}$. Then by using Lemma 1, we know that this $\epsilon$-accurate stationary point is also an $\epsilon$-accurate Pareto stationary point of normalized functions $\{\tilde{l}_{i}(x)\}$ satisfying
$\min_{w\in\mathcal{W}}\|G(x^{T})w\|^{2}=\mathcal{O}(\epsilon).$
The proof of the first half of our theorem is complete.
Then for the second half, since we have built the connection between the stationarity of $\mathcal{BP}_{\lambda}$ and Pareto stationarity, we prove that the single-loop Algorithm 1 achieves an $\epsilon$-accurate stationary point of $\mathcal{BP}_{\lambda}$. Recall that
$\|\nabla f(W^{t},$ $x^{t})+\lambda(\nabla g(W^{t},x^{t})-\nabla g(W^{t},x_{t}^{*}))\|^{2}$
$\overset{(i)}{\leq}$ $2\|\nabla f(W^{t},x^{t})+\lambda\nabla g(W^{t},x^{t})\|^{2}+2\lambda^{2}\|\nabla g(W^{t},x_{t}^{*})\|^{2}$
=$\overset{(ii)}{=}$ $2\|\nabla f(W^{t},x^{t})+\lambda\nabla g(W^{t},x^{t})\|^{2}+2\lambda^{2}\|\nabla_{W}g(W^{t},x_{t}^{*})\|^{2}$
$\overset{(iii)}{\leq}$ $2\|\nabla f(W^{t},x^{t})+\lambda\nabla g(W^{t},x^{t})\|^{2}+4\lambda^{2}\|\nabla_{W}g(W^{t},x^{*}_{t})-\nabla_{W}g(W^{t},z^{t}_{N})\|^{2}+4\lambda^{2}\|\nabla_{W}g(W^{t},z^{t}_{N})\|^{2},$ (10)
where $(i)$ and $(iii)$ both follow from Young’s inequality, and $(ii)$ follows from $\nabla_{x}g(W^{t},x_{t}^{*})=0$. Besides, recall that $z_{N}^{t}$ is the intermediate output of the subloop in Algorithm 2. We next provide the upper bounds of the above three terms on the right-hand side (RHS). For the first term, we utilize the smoothness of $\Phi(\theta^{t})=f(\theta^{t})+\lambda g(\theta^{t})$ where $L_{\Phi}=L_{f}+\lambda L_{g}$ and $\theta^{t}=(W^{t},x^{t})$.
$\Phi(\theta^{t+1})\leq$ $\Phi(\theta^{t})+\langle\nabla\Phi(\theta^{t}),\theta^{t+1}-\theta^{t}\rangle+\frac{L_{\Phi}}{2}\|\theta^{t+1}-\theta^{t}\|^{2}$
$\overset{(i)}{\leq}$ $\Phi(\theta^{t})-\frac{\alpha}{2}\|\nabla\Phi(\theta^{t})\|^{2},$
where $(i)$ follows from $\alpha\leq\frac{1}{L_{\Phi}}=\mathcal{O}(\lambda^{-1})$. Thus, we can obtain
$\|\nabla\Phi(\theta^{t})\|^{2}\leq\frac{2}{\alpha}(\Phi(\theta^{t})-\Phi(\theta^{t+1})).$ (11)
Then for the second term on the RHS in appendix C, we follow the same step in the proof of Theorem 3 in Shen & Chen 2023 and obtain
$4\lambda^{2}\|\nabla_{W}$ $g(W^{t},x^{*}_{t})-\nabla_{W}g(W^{t},z_{N}^{t})\|^{2}$
$\leq$ $4\lambda^{2}L_{g}^{2}\mu\Big{(}1-\frac{\beta}{2\mu}\Big{)}^{N}(g(W^{t},x^{t})-g(W^{t},x^{*}_{t}))$
$\overset{(i)}{\leq}$ $4\lambda^{2}L_{g}^{2}\Big{(}1-\frac{\beta}{2\mu}\Big{)}^{N}\|\nabla_{x}g(W^{t},x^{t})\|^{2}$
$=$ $4\lambda^{2}L_{g}^{2}\Big{(}1-\frac{\beta}{2\mu}\Big{)}^{N}\Big{\|}\frac{x^{t+1}-x^{t}+\alpha\nabla_{x}f(W^{t},x^{t})}{\alpha\lambda}\Big{\|}^{2}$
$\overset{(ii)}{\leq}$ $8\lambda^{2}L_{g}^{2}\Big{(}1-\frac{\beta}{2\mu}\Big{)}^{N}\Big{(}\frac{\|\theta^{t+1}-\theta^{t}\|^{2}}{\alpha^{2}\lambda^{2}}+\frac{L^{2}}{\lambda^{2}}\Big{)}$
$\overset{(iii)}{\leq}$ $\frac{1}{2\alpha^{2}}\|\theta^{t+1}-\theta^{t}\|^{2}+\frac{2L^{2}L_{g}^{2}}{\alpha^{2}t^{2}}$
$=$ $\frac{1}{2}\|\nabla\Phi(\theta^{t})\|^{2}+\frac{2L^{2}L_{g}^{2}}{\alpha^{2}t^{2}},$ (12)
where $(i)$ follows from the PL condition, $(ii)$ follows from Young’s inequality and 1, and $(iii)$ follows from the selection on $N\geq\max\{-\log_{c_{\beta}}(16L_{g}^{2}),-2\log_{c_{\beta}}(2\alpha t)\}$ with $c_{\beta}=1-\frac{\beta}{2\mu}$. Lastly, for the last term at the RHS in appendix C, we have,
$4\lambda^{2}\|\nabla_{W}g(W^{t},z^{t}_{N})\|^{2}=\mathcal{O}(\lambda^{2}\epsilon^{2}),$ (13)
where this inequality follows from our experimental observation. Furthermore, substituting eq. 11, and appendix C into appendix C yields
$\|\nabla f(W^{t},x^{t})$ $+\lambda(\nabla g(W^{t},x^{t})-\nabla g(W^{t},x^{*}_{t}))\|^{2}$
$\leq$ $\frac{5}{2}\|\nabla\Phi(\theta^{t})\|^{2}+\frac{2L^{2}L_{g}^{2}}{\alpha^{2}t^{2}}+4\lambda^{2}\|\nabla_{W}g(W^{t},z^{t}_{N})\|^{2}$
$\leq$ $\frac{5}{\alpha}(\Phi(\theta^{t})-\Phi(\theta^{t+1}))+\frac{2L^{2}L_{g}^{2}}{\alpha^{2}t^{2}}+4\lambda^{2}\|\nabla_{W}g(W^{t},z^{t}_{N})\|^{2}.$ (14)
Therefore, telescoping the above inequality yields,
$\frac{1}{T}\sum_{t=0}^{T-1}$ $\|\nabla f(W^{t},x^{t})+\lambda(\nabla g(W^{t},x^{t})-\nabla g(W^{t},x^{*}_{t}))\|^{2}$
$=$ $\mathcal{O}(\frac{\lambda}{\alpha T}+\frac{1}{\alpha^{2}T}+\lambda^{2}\epsilon^{2}).$
According to the parameter selection that $\lambda=\mathcal{O}(\epsilon^{-\frac{1}{2}})$, $\alpha=\mathcal{O}(\epsilon^{\frac{1}{2}})$, and $T=\Omega(\epsilon^{-2})$, we can obtain
$\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla f(W^{t},x^{t})+\lambda(\nabla g(W^{t},x^{t})-\nabla g(W^{t},x^{*}_{t}))\|^{2}=\mathcal{O}(\epsilon).$
Therefore, Algorithm 1 can achieve a stationary point of $\mathcal{BP}_{\lambda}$ with $\mathcal{O}(\epsilon^{-2})$ iterations. If this stationary point is a local/global solution to $\mathcal{BP}_{\lambda}$, it is also a solution to $\mathcal{BP}_{\epsilon}$ according to Proposition 2 in Shen & Chen 2023. Then by using Lemma 1, we know this stationary point is also an $\epsilon$-accurate Pareto stationary point of the original loss functions. The proof is complete. ∎